You decide to finance a laptop. What will your monthly payments be if you finance $2500 at 7% for 3 years? What will the acquiaition be?
Answer #1

Solution :-

Annual rate = 7%

So monthly rate = 7%/12 = 0.5833%

Payment in a year = 12

No. Of years = 3

Therefore total payments = 12*3 = 36 payments

Principal amount of laptop = $2500

Therefore monthly payments = $2500 / PVAF(0.5833%, 36)

= $2500/32.3865 = $77.192

Acquiation = $77.192*36 - $2500 = $278.936